3x^2+3x-3 and need to plug in f(3x)
Found 2 solutions by Boreal, fractalier:
Answer by Boreal(15235)

everywhere there is an x, put 3x
3(3x)^2+3(3x)+3=
27x^2+9x+3, remembering that (3x)^2=3^2*x^2

Answer by fractalier(6550)

Same as the last one...just plug 3x in for every x you see in the original...
f(3x) = 3(3x)^2 + 3(3x) - 3 = 27x^2 + 9x - 3